# Polylog-time- and constant-space-overhead fault-tolerant quantum computation with quantum low-density parity-check codes

### v1

@@ Abstract
thanks: This paper is a full paper of the work that has been presented in Refs. [1, 2].
@@ III Setting of fault-tolerant quantum computation
Let $C^{\mathrm{FT}}$ be a fault-tolerant circuit. For simplicity, the set of all the physical locations of the fault-tolerant circuit is denoted also by $C^{\mathrm{FT}}$. A set of faulty locations is given by a random variable $F\subseteq C^{\mathrm{FT}}$ of locations of the circuit, where any Pauli errors may occur. We say that $C^{\mathrm{FT}}$ experiences a local stochastic Pauli error model with error parameters $\{p_{i}\}_{i\in C^{\mathrm{FT}}}$ if the following conditions hold.
@@
    Each physical location $i\in C^{\mathrm{FT}}$ has an error parameter $p_{i}\in(0,1]$ such that for all $A\subseteq L$, the probability that $F$ contains $A$ satisfies
@@ V.1 Conditions of fault-tolerant gadgets for quantum LDPC codes
In the following, we define two decoding algorithms that quantum LDPC codes used in our protocol should have: the single-shot decoding algorithm with thresholds for reducing residual errors in Def. 5 and the decoding algorithm with threshold for correcting errors in Def. 6. These algorithms are designed to correct local stochastic Pauli errors on physical qubits and syndrome bits with high probability if an error parameter is below a certain threshold value. The family of quantum expander codes with $D=\Theta(\sqrt{N})$ is the only known family of non-vanishing-rate quantum LDPC codes satisfying the requirements in Defs. 5 and 6. Specifically, the small-set-flip decoding algorithm for the quantum expander codes can work as both the single-shot decoding algorithm for reducing errors with thresholds in Def. 5 and the decoding algorithm for correcting errors in Def. 6 [29, 28] (in particular, see Theorems 5.30 and 5.36 in [29]) More recently, Ref. [68] showed that the family of quantum Tanner codes with $D=\Theta(N)$ also has a single-shot decoding algorithm in the sense of Def. 93, but there is no proof that bounds the error parameter of a residual error under the local stochastic error model as in (97). Although it is conceivable that quantum Tanner codes could also achieve polylogarithmic time and constant space overhead FTQC based on our protocol, we leave its rigorous proof open. Note that our analysis on the threshold theorem and the space and time overheads in Sec. VI are independent of any other properties of quantum expander codes and thus will work for the quantum Tanner codes if statements on the decoding algorithms that can be used in place of Defs. 93 and (97) are established.
@@
$\mathbb{P}[m(e_{P},\Delta_{P})=0\text{~{}and~{}}\mathrm{supp}(e^{\prime}_{P})\supseteq S^{\prime}]\leq\quantity(p_{1}^{\mathrm{th}})^{c(T)},$ (97)
@@ V.2.8 Error-correction gadget
where $c(T)$ is the error suppression parameter, which is a monotonically increasing function with respect to $T$ and $p_{\mathrm{dec}}^{\mathrm{th}}\coloneqq\min\{p_{1}^{\mathrm{th}},p_{2}^{\mathrm{th}}\}>0$ where $p_{1}^{\mathrm{th}}$ and $p_{2}^{\mathrm{th}}$ are the threshold constants of the decoding algorithm in Defs. 5 and 6. Here, $\lambda$ is a positive constant satisfying, with some constant $d_{0}\geq d^{\mathrm{synd}}$,
$\lambda<\quantity(\frac{(1-p_{\mathrm{dec}}^{\mathrm{th}})p_{\mathrm{dec}}^{\mathrm{th}}}{2^{2d_{0}}})^{2^{2d_{0}}}.$ (133)
With this choice, $T$ does not depend on $N$. The width of the gadget is
@@
where $M=N-K$ is the number of auxiliary physical qubits for measuring stabilizer generators.
Therefore, the EC gadget is fault-tolerant in the sense of Def. 10.
@@ VI Threshold theorem and overhead analysis
(Threshold theorem for polylog-time- and constant-space-overhead fault-tolerant quantum computation). Let $\quantity{C^{\mathrm{org}}_{n}}_{n}$ be a sequence of original circuits, where each circuit $C_{n}^{\mathrm{org}}$ is specified by an integer $n$ and has width $W(n)\rightarrow\infty$ and depth $D(n)=O(\mathrm{poly}(W(n)))$ as $n\to\infty$. Suppose that an original circuit $C^{\mathrm{org}}_{n}$ is compiled into a fault-tolerant circuit $C^{\mathrm{FT}}_{n}$ as described in Sec. IV.1, and $C^{\mathrm{FT}}_{n}$ experiences a local stochastic Pauli error model with error parameters
$p_{i}\leq p_{\mathrm{loc}}\text{ ~{}for all locations ~{}}i\in C_{n}^{\mathrm{FT}},$ (135)
satisfying Assumption 2.
@@
Due to (136), there exist positive constants $c_{\max}$ and $c_{\min}$ ($c_{\max}\geq c_{\min}>0$) such that for sufficiently large $i$
$c_{\min}D_{i}\leq N_{i}^{\gamma}\leq c_{\max}D_{i}.$ (141)
For any positive constant $c_{\mathrm{dist}}>0$, we define the lower bound $D_{\min}$ for the code distance as
$D_{{\min}}\coloneqq c_{\mathrm{dist}}\log^{\alpha\gamma}\quantity(\frac{|C_{n}^{\mathrm{org}}|}{\varepsilon}).$ (142)
@@
$i_{\min}\coloneqq\min\{i\colon\mathcal{Q}_{i}\text{~{}with~{}}D_{i}\geq D_{\mathrm{min}}\}.$ (143)
@@
Due to (137), for sufficiently large $n$ satisfying (141), we have
$c_{\min}D_{i_{\min}}^{\gamma}\leq N_{i_{\min}}=N(n,\varepsilon)$ $\leq N_{i_{\min}-1}+\quantity(N_{i_{\min}-1})^{\beta}$ (145)
@@
$N(n,\varepsilon)=\Theta\quantity(\log^{\alpha}\quantity(\frac{|C_{n}^{\mathrm{org}}|}{\varepsilon})),$ (149)
@@
$\lambda<\quantity(\frac{p_{\mathrm{dec}}^{\mathrm{th}}}{\omega 2^{2d_{0}}})^{2^{2d_{0}}},$ (161)
where
$\omega\coloneqq(1-p_{\mathrm{dec}}^{\mathrm{th}})^{-1}.$ (162)
@@ VI.2 Partial circuit reduction
A pair $\quantity(C,\{p_{i}\}_{i\in C})$, where $C$ is partitioned into two disjoint subsets $C_{1}$ and $C_{2}$, i.e., $C=C_{1}\cup C_{2}$ and $C_{1}\cap C_{2}=\varnothing$, such that
$\forall i\in C_{1},\quad p_{i}$ $\leq p^{(1)}\in(0,1],$ (235)
$\forall i\in C_{2},\quad p_{i}$ $\leq p^{(2)}\in(0,1],$ (236)
is reducible (i.e., $\delta$-reducible with $\delta=0$) to another pair $\quantity(C^{\prime},\{q_{j}\}_{j\in C^{\prime}})$ if there exists a map $\Gamma\colon 2^{C}\to 2^{C^{\prime}}$ and positive integers $c_{\rm{fw}}$, $c_{\rm{bw}}^{(1)}$, and $c_{\rm{bw}}^{(2)}$ satisfying the following conditions.
@@
    $\Gamma(A)=\bigcup_{i\in A}\Gamma(\{i\})\text{~{}and~{}}\Gamma(\varnothing)=\varnothing$.
@@
    |$|\{i\in C_{1}\colon j\in\Gamma(\{i\})\}|\leq c_{\rm{bw}}^{(1)}\text{~{}and~{}}\{i\in C_{2}\colon j\in\Gamma(\{i\})\}|\leq c_{\rm{bw}}^{(2)}\text{~{}for all~{}}j\in C^{\prime}$.
@@
    $q_{j}\geq 2^{c_{\mathrm{bw}}^{(2)}}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{(1)}}\frac{\quantity(p^{(2)})^{1/c_{\mathrm{fw}}}}{1-p^{(2)}}\text{~{}for all~{}}j\in C^{\prime}$.
@@
    For any set $A\subseteq C$ of faulty locations in $C$ and any assignment of Pauli errors in $A$, there exists an assignment of Pauli errors in the set $\Gamma(A)\subseteq C^{\prime}$ in $C^{\prime}$ such that $C$ with the errors in $A$ and $C^{\prime}$ with the errors in $\Gamma(A)$ are equivalent.
@@
$\mathbb{P}[F\cap C_{1}\supseteq A_{1}\text{~{}and~{}}F\cap C_{2}\supseteq A_{2}]\leq\quantity(p^{(1)})^{|A_{1}|}\quantity(p^{(2)})^{|A_{2}|}.$ (241)
Our goal is to show that
@@
for $|A^{\prime}|\geq 1$.
If $\Gamma(F)\supseteq A^{\prime}$, by condition 1,
@@
where $\Lambda(A^{\prime})\coloneqq\Lambda_{1}(A^{\prime})\cup\Lambda_{2}(A^{\prime})$. Thus, we have
$|A^{\prime}|\leq\sum_{i\in F\cap\Lambda(A^{\prime})}|\Gamma(\{i\})|\leq c_{\mathrm{fw}}|F\cap\Lambda(A^{\prime})|$ (246)
using condition 2. Hence, it holds that
$\mathbb{P}[\Gamma(F)\supseteq A^{\prime}]$ $\leq\mathbb{P}\left[|F\cap\Lambda(A^{\prime})|\geq\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right]$ (247)
$=\sum_{H\subseteq\Lambda(A^{\prime}):|H|\geq|A^{\prime}|/c_{\mathrm{fw}}}\mathbb{P}[F\cap\Lambda(A^{\prime})=H]$ (248)
$\leq\sum_{H\subseteq\Lambda(A^{\prime}):|H|\geq|A^{\prime}|/c_{\mathrm{fw}}}\mathbb{P}[F\supseteq H].$ (249)
To bound this probability, we consider all subsets $H\subseteq\Lambda(A^{\prime})$ of size $r\geq\lceil|A^{\prime}|/c_{\mathrm{fw}}\rceil$. Then,
@@
$\mathbb{P}\left[|F\cap\Lambda(A^{\prime})|\geq\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right]=\sum_{r=\left\lceil\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right\rceil}^{|\Lambda(A^{\prime})|}\sum_{\begin{subarray}{c}H\subseteq\Delta(A^{\prime})\\ |H|=r\end{subarray}}\mathbb{P}[F\supseteq H].$ (250)
For each such $H$, we partition it into $H_{1}=H\cap C_{1}$ and $H_{2}=H\cap C_{2}$, with sizes $k=|H_{1}|$ and $r-k=|H_{2}|$. Using the assumption (241) on $F$, we have
$\mathbb{P}[F\supseteq H]\leq\prod_{i\in H_{1}}p_{i}\prod_{j\in H_{2}}p_{j}\leq\quantity(p^{(1)})^{k}\quantity(p^{(2)})^{r-k}.$ (251)
The number of choices of $H_{1}$ and $H_{2}$ is bounded by
$\binom{|\Lambda_{1}(A^{\prime})|}{k}\binom{|\Lambda_{2}(A^{\prime})|}{r-k}\leq\binom{c_{\mathrm{bw}}^{(1)}|A^{\prime}|}{k}\binom{c_{\mathrm{bw}}^{(2)}|A^{\prime}|}{r-k}.$ (252)
Thus,

### v2

@@ III Setting of fault-tolerant quantum computation
Let $C^{\mathrm{FT}}$ be a fault-tolerant circuit. For simplicity, the set of all the physical locations of the fault-tolerant circuit is also denoted by $C^{\mathrm{FT}}$. A set of faulty locations is given by a random variable $F\subseteq C^{\mathrm{FT}}$ of locations of the circuit, where any Pauli errors may occur. We say that $C^{\mathrm{FT}}$ experiences a local stochastic Pauli error model with error parameters $\{p_{i}\}_{i\in C^{\mathrm{FT}}}$ if the following conditions hold.
@@
    Each physical location $i\in C^{\mathrm{FT}}$ has an error parameter $p_{i}\in(0,1]$ such that for all $A\subseteq C^{\mathrm{FT}}$, the probability that $F$ contains $A$ satisfies
@@ V.1 Conditions of fault-tolerant gadgets for quantum LDPC codes
In the following, we define two decoding algorithms that quantum LDPC codes used in our protocol should have: the single-shot decoding algorithm with thresholds for reducing residual errors in Def. 5 and the decoding algorithm with threshold for correcting errors in Def. 6. These algorithms are designed to correct local stochastic Pauli errors on physical qubits and syndrome bits with high probability if an error parameter is below a certain threshold value. The family of quantum expander codes with $D=\Theta(\sqrt{N})$ is the only known family of non-vanishing-rate quantum LDPC codes satisfying the requirements in Defs. 5 and 6. Specifically, the small-set-flip decoding algorithm for the quantum expander codes can work as both the single-shot decoding algorithm for reducing errors with thresholds in Def. 5 and the decoding algorithm for correcting errors in Def. 6 [27, 26] (in particular, see Theorems 5.30 and 5.36 in [27]) More recently, Ref. [66] showed that the family of quantum Tanner codes with $D=\Theta(N)$ also has a single-shot decoding algorithm in the sense of Def. 93, but there is no proof that bounds the error parameter of a residual error under the local stochastic error model as in (97). Building on this direction, Ref. [73] extends the analysis to provide an error bound for error correction with quantum Tanner codes within the EC gadget. However, Ref. [73] does not provide a complete error analysis for all gadgets but focuses on analyzing specific parts of the fault-tolerant protocol. In contrast, our contribution lies in performing a comprehensive analysis for all gadgets, thereby completing the proof of the threshold theorem.
@@
$\mathbb{P}[m(e_{P},\Delta_{P})=0\text{~{}and~{}}\mathrm{supp}(e^{\prime}_{P})\supseteq S^{\prime}]\leq\quantity(p_{1}^{\mathrm{th}})^{c(T)|S^{\prime}|},$ (97)
@@ V.2.8 Error-correction gadget
where $c(T)$ is the error suppression parameter, which is a monotonically increasing function with respect to $T$ and $p_{\mathrm{dec}}^{\mathrm{th}}\coloneqq\min\{p_{1}^{\mathrm{th}},p_{2}^{\mathrm{th}}\}>0$ where $p_{1}^{\mathrm{th}}$ and $p_{2}^{\mathrm{th}}$ are the threshold constants of the decoding algorithm in Defs. 5 and 6. With this choice, $T$ does not depend on $N$. Here, $\lambda$ is a positive constant satisfying, with some constant $d_{0}\geq d^{\mathrm{synd}}$,
$\lambda<\quantity(\frac{p_{\mathrm{dec}}^{\mathrm{th}}}{2^{2d_{0}}})^{2^{2d_{0}}}.$ (133)
The width of the gadget is
@@
where $M=N-K$ is the number of auxiliary physical qubits for measuring stabilizer generators. Therefore, the EC gadget is fault-tolerant in the sense of Def. 10.
@@ VI Threshold theorem and overhead analysis
(Threshold theorem for polylog-time- and constant-space-overhead fault-tolerant quantum computation). Let $\quantity{C^{\mathrm{org}}_{n}}_{n}$ be a sequence of original circuits, where each circuit $C_{n}^{\mathrm{org}}$ is specified by an integer $n$ and has width $W(n)\rightarrow\infty$ and depth $D(n)=O(\mathrm{poly}(W(n)))$ as $n\to\infty$. Suppose that an original circuit $C^{\mathrm{org}}_{n}$ is compiled into a fault-tolerant circuit $C^{\mathrm{FT}}_{n}$ as described in Sec. IV.1, and $C^{\mathrm{FT}}_{n}$ experiences a local stochastic Pauli error model with error parameters $p_{i}\leq p_{\mathrm{loc}}\text{ ~{}for all locations ~{}}i\in C_{n}^{\mathrm{FT}}$ satisfying Assumption 2.
@@
For any positive constant $c_{N}>0$, we define the lower bound $N_{\min}$ for the code distance as
$N_{{\min}}\coloneqq c_{N}\log^{\alpha}\quantity(\frac{|C_{n}^{\mathrm{org}}|}{\varepsilon}).$ (140)
@@
$i_{\min}\coloneqq\min\{i\colon\mathcal{Q}_{i}\text{~{}with~{}}N_{i}\geq N_{\mathrm{min}}\}.$ (141)
@@
Due to (136), we have
$N_{\min}\leq N_{i_{\min}}=N(n,\varepsilon)$ $\leq N_{i_{\min}-1}+\quantity(N_{i_{\min}-1})^{\beta}$ (143)
@@
$N(n,\varepsilon)=\Theta\quantity(\log^{\alpha}\quantity(\frac{|C_{n}^{\mathrm{org}}|}{\varepsilon})).$ (146)
Due to (142), for sufficiently large $n$, we have
@@
$\lambda<\quantity(\frac{p_{\mathrm{dec}}^{\mathrm{th}}}{2^{2d_{0}}})^{2^{2d_{0}}},$ (158)
@@ VI.2 Partial circuit reduction
Consider a pair $\quantity(C,\{p_{i}\}_{i\in C})$ for which $C$ is partitioned into two disjoint subsets $C_{1}$ and $C_{2}$, and
$\forall i\in C_{1},\quad p_{i}$ $\leq p^{(1)}\in[0,1],$ (231)
$\forall i\in C_{2},\quad p_{i}$ $\leq p^{(2)}\in(0,1]$ (232)
holds. The pair $\quantity(C,\{p_{i}\}_{i\in C})$ is reducible (i.e., $\delta$-reducible with $\delta=0$) to another pair $\quantity(C^{\prime},\{q_{j}\}_{j\in C^{\prime}})$ if there exists a map $\Gamma\colon 2^{C}\to 2^{C^{\prime}}$ and positive integers $c_{\rm{fw}}$, $c_{\rm{bw}}^{(1)}$, and $c_{\rm{bw}}^{(2)}$ satisfying the following conditions.
@@
    It holds that
    $\Gamma(A)$ $=\bigcup_{i\in A}\Gamma(\{i\}),$ (233)
    $\Gamma(\varnothing)$ ==$\varnothing$.= ∅ . (234)
@@
    For all $j\in C^{\prime}$,
    |$|\{i\in C_{1}\colon j\in\Gamma(\{i\})\}|$ $\leq c_{\rm{bw}}^{(1)},$ (236)
    |$|\{i\in C_{2}\colon j\in\Gamma(\{i\})\}|$ $\leq c_{\rm{bw}}^{(2)}.$ (237)
@@
    Define
    $C_{\mathrm{faulty}}^{\prime}\coloneqq\bigcup_{i\in C:p_{i}\neq 0}\Gamma(\{i\}).$ (238)
    For all $j\in C^{\prime}_{\mathrm{faulty}}$,
    $q_{j}\geq 2^{c_{\mathrm{bw}}^{(2)}}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\mathrm{bw}}^{(1)}}\quantity(p^{(2)})^{1/c_{\mathrm{fw}}}$
    $+2^{c_{\mathrm{bw}}^{(1)}}\quantity(p^{(1)})^{1/c_{\mathrm{fw}}};$ (239)
    and for all $j\in C^{\prime}\setminus C^{\prime}_{\mathrm{faulty}}$,
    $q_{j}=0.$ (240)
@@
    For any set $A\subseteq C$ of faulty locations in $C$ and any assignment of Pauli errors in $A$, there exists an assignment of Pauli errors in the set $\Gamma(A)\subseteq C^{\prime}$ such that $C$ with the errors in $A$ and $C^{\prime}$ with the errors in $\Gamma(A)$ are equivalent.
@@
$\mathbb{P}[F\cap C_{1}\supseteq A_{1}\text{~{}and~{}}F\cap C_{2}\supseteq A_{2}]\leq\quantity(p^{(1)})^{|A_{1}|}\quantity(p^{(2)})^{|A_{2}|},$ (245)
where $10^{0}=$ 1 when $p^{(1)}=0$ and $|A_{1}|=0$. Our goal is to show that
@@
for $|A^{\prime}|\geq 1$. In particular, if $A^{\prime}\not\subseteq C_{\mathrm{faulty}}^{\prime}$, then $\mathbb{P}[\Gamma(F)\supseteq A^{\prime}]=0$; thus, the goal reduces to showing
[…]
for $|A^{\prime}|\geq 1$ and $q_{j}$ satisfying (4). In the following analysis, we consider $A^{\prime}\subseteq C_{\mathrm{faulty}}^{\prime}$.
If $\Gamma(F)\supseteq A^{\prime}$, defining $\Lambda(A^{\prime})\coloneqq\Lambda_{1}(A^{\prime})\cup\Lambda_{2}(A^{\prime})$, we obtain from condition 1
@@
Thus, using condition 2, we have
$|A^{\prime}|\leq\sum_{i\in F\cap\Lambda(A^{\prime})}|\Gamma(\{i\})|\leq c_{\mathrm{fw}}|F\cap\Lambda(A^{\prime})|.$ (251)
We write
@@
$k_{0}\coloneqq\left\lceil\frac{|A^{\prime}|}{c_{\mathrm{fw}}}\right\rceil\geq\frac{|A^{\prime}|}{c_{\mathrm{fw}}}.$ (252)
Then, it holds that
$\mathbb{P}[\Gamma(F)\supseteq A^{\prime}]$ $\leq\mathbb{P}\left[c_{\mathrm{fw}}|F\cap\Lambda(A^{\prime})|\geq|A^{\prime}|\right]$ (253)
$=\mathbb{P}\left[|F\cap\Lambda(A^{\prime})|\geq k_{0}\right]$ (254)
To bound this probability, we consider two cases based on the size of $|F\cap\Lambda_{1}(A^{\prime})|$, i.e.,
$\mathbb{P}\left[|F\cap\Lambda(A^{\prime})|\geq k_{0}\right]$
$\leq\mathbb{P}[|F\cap\Lambda_{1}(A^{\prime})|\geq k_{0}]+$
$~{}~{}~{}~{}\mathbb{P}[|F\cap\Lambda(A^{\prime})|\geq k_{0}\text{~{}and~{}}|F\cap\Lambda_{1}(A^{\prime})|\leq k_{0}].$ (255)
In the following, we will bound each term on the right-hand side.
For the first term, it hold that
$\mathbb{P}[|F\cap\Lambda_{1}(A^{\prime})|\geq k_{0}]$
$=\sum_{\begin{subarray}{c}H\subseteq\Lambda_{1}(A^{\prime}):\\ |H|\geq k_{0}\end{subarray}}\mathbb{P}[F\cap\Lambda_{1}(A^{\prime})=H]$ (256)
$\leq\sum_{\begin{subarray}{c}H\subseteq\Lambda_{1}(A^{\prime}):\\ |H|\geq k_{0}\end{subarray}}\quantity(p^{(1)})^{k_{0}}$ (257)
$\leq\sum_{H\subseteq\Lambda_{1}(A^{\prime})}\quantity(p^{(1)})^{k_{0}}$ (258)
$=2^{|\Lambda_{1}(A^{\prime})|}\quantity(p^{(1)})^{k_{0}},$ (259)
where we used (245). Due to (243), this becomes
$\mathbb{P}[|F\cap\Lambda_{1}(A^{\prime})|\geq k_{0}]\leq\left(2^{c_{\mathrm{bw}}^{(1)}}\quantity(p^{(1)})^{1/c_{\mathrm{fw}}}\right)^{|A^{\prime}|},$ (260)
where we used (252) and $p^{(1)}\leq 1$.
For the second term, we have
$\mathbb{P}[|F\cap\Lambda(A^{\prime})|\geq k_{0}\text{~{}and~{}}|F\cap\Lambda_{1}(A^{\prime})|\leq k_{0}]$
$=\sum_{\begin{subarray}{c}H\subseteq\Lambda(A^{\prime}):\\ |H|\geq k_{0}\\ |H\cap\Lambda_{1}(A^{\prime})|\leq k_{0}\end{subarray}}\mathbb{P}[F\cap\Lambda(A^{\prime})=H].$ (261)
We further decompose $H$ on the right-hand side into disjoint sets $H_{1}$ and $H_{2}$, i.e., $H_{1}=H\cap\Lambda_{1}(A^{\prime})$ and $H_{2}=H\cap\Lambda_{2}(A^{\prime})$, so $H=H_{1}\cup H_{2}$. Then we have
$\mathbb{P}[|F\cap\Lambda(A^{\prime})|\geq k_{0}\text{~{}and~{}}|F\cap\Lambda_{1}(A^{\prime})|\leq k_{0}]$
$=\sum_{\begin{subarray}{c}H_{1}\subseteq\Lambda_{1}(A^{\prime}):\\ |H_{1}|\leq k_{0}\end{subarray}}\sum_{\begin{subarray}{c}H_{2}\subseteq\Lambda_{2}(A^{\prime}):\\ |H_{1}|+|H_{2}|\geq k_{0}\end{subarray}}\mathbb{P}[F\cap\Lambda(A^{\prime})=H_{1}\cup H_{2}]$ (262)
$=\sum_{r=0}^{k_{0}}\sum_{\begin{subarray}{c}H_{1}\subseteq\Lambda_{1}(A^{\prime}):\\ |H_{1}|=r\end{subarray}}\sum_{\begin{subarray}{c}H_{2}\subseteq\Lambda_{2}(A^{\prime}):\\ |H_{2}|\geq k_{0}-r\end{subarray}}\mathbb{P}[F\cap\Lambda(A^{\prime})=H_{1}\cup H_{2}]$ (263)
$\leq\sum_{r=0}^{k_{0}}\sum_{\begin{subarray}{c}H_{1}\subseteq\Lambda_{1}(A^{\prime}):\\ |H_{1}|=r\end{subarray}}\sum_{\begin{subarray}{c}H_{2}\subseteq\Lambda_{2}(A^{\prime}):\\ |H_{2}|\geq k_{0}-r\end{subarray}}\quantity(p^{(1)})^{r}\quantity(p^{(2)})^{k_{0}-r}$ (264)
$\leq\sum_{r=0}^{k_{0}}\sum_{\begin{subarray}{c}H_{1}\subseteq\Lambda_{1}(A^{\prime}):\\ |H_{1}|=r\end{subarray}}\sum_{H_{2}\subseteq\Lambda_{2}(A^{\prime})}\quantity(p^{(1)})^{r}\quantity(p^{(2)})^{k_{0}-r}$ (265)
$=\sum_{r=0}^{k_{0}}\binom{|\Lambda_{1}(A^{\prime})|}{r}2^{|\Lambda_{2}(A^{\prime})|}\quantity(p^{(1)})^{r}\quantity(p^{(2)})^{k_{0}-r}$ (266)
$\leq 2^{|\Lambda_{2}(A^{\prime})|}\sum_{r=0}^{|\Lambda_{1}(A^{\prime})|}\binom{|\Lambda_{1}(A^{\prime})|}{r}\quantity(p^{(1)})^{r}\quantity(p^{(2)})^{k_{0}-r}$ (267)
$=2^{|\Lambda_{2}(A^{\prime})|}\quantity(p^{(2)})^{k_{0}-|\Lambda_{1}(A^{\prime})|}\quantity(p^{(1)}+p^{(2)})^{|\Lambda_{1}(A^{\prime})|}$ (268)
$=2^{|\Lambda_{2}(A^{\prime})|}\quantity(p^{(2)})^{k_{0}}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{|\Lambda_{1}(A^{\prime})|}.$ (269)
where we used (245) for the inequality (264), and (269) follows from $p^{(2)}>0$. Using the bounds (243), (244), and (252), and $p^{(2)}\leq 1$, we have
$\mathbb{P}[|F\cap\Lambda(A^{\prime})|\geq k_{0}\text{ and }|F\cap\Lambda_{1}(A^{\prime})|\leq k_{0}]$ (270)
$\leq\left(2^{c_{\mathrm{bw}}^{(2)}}\left(1+\frac{p^{(1)}}{p^{(2)}}\right)^{c_{\text{bw}}^{(1)}}\quantity(p^{(2)})^{1/c_{\text{fw}}}\right)^{|A^{\prime}|}.$ (271)
Adding the bounds from the two terms, and noting that $x^{n}+y^{n}\leq(x+y)^{n}$ for $x,y\geq 0$ and $n\geq 1$, we obtain